OpenStudy (anonymous):

Find the function F(x), given that F'(x)=(6/x^2)- (10/x^6) and F(1)=0

OpenStudy (irishboy123):

\( F'(x)=\dfrac{6}{x^2} - \dfrac{10}{x^6} \) \( F(x)=\int ~ \dfrac{6}{x^2} - \dfrac{10}{x^6} ~dx \) \(F(1) = 0\)

OpenStudy (irishboy123):

nuff said ?!?!

OpenStudy (anonymous):

So F(x)= (2/x^5) + (6/x) +C ??? @IrishBoy123

OpenStudy (irishboy123):

\(F(x)=\int ~ \dfrac{6}{x^2} - \dfrac{10}{x^6} ~dx\) \(= \dfrac{2}{x^5} - \dfrac{6}{x} + C\) \(F(1) = 0 \implies 0 = 2 - 6 + C \implies C = 4\) \(F(x) = \dfrac{2}{x^5} - \dfrac{6}{x} + 4\)

OpenStudy (anonymous):

Oh I see how that works now! Thank you!